# Quantum teleportation implies symmetry-protected topological order

### v2

@@ Appendix A Proof of Proposition 1
An analogous procedure applies to the $Z$-type logical operator for the $n$th logical qubit. That operator evolves in the Heisenberg picture according to
$Z^{(n)}_{{\rm L}}(T)\,=\,Z_{f_{n}}(T)\,=\,\mathcal{W}^{\dagger}\,Z_{f_{n}}\,\mathcal{W}\,,~{}~{}$ (A.92)
@@
in analogy to Eq. A.87, and implied by Eq. 7. We then define another stabilizer $\mathsf{S}_{n,z}\equiv Z_{i_{n}}\,\mathcal{W}^{\dagger}Z_{f_{n}}\mathcal{W}$, implying that $\mathsf{S}_{n,z}$ is also an element of $\operatorname*{stab}\left(*\right){\lvert\Psi_{0}\rangle}$ (3). Applying the logical $Z$ operator (A.92) to $\lvert\Psi_{0}\rangle$ (6) reveals that
$Z^{(n)}_{{\rm L}}(T)=\mathcal{W}^{\dagger}\,Z_{f_{n}}\,\mathcal{W}=Z_{i_{n}}\,\mathsf{S}_{n,z}\,,~{}~{}$ (A.93)
@@
is the $Z$-type logical operator for the $n$th logical qubit acting on the initial state $\lvert\Psi_{0}\rangle$ (6), in analogy to Eq. A.91.
@@
Thus, the fact that $\mathcal{W}$ teleports $k$ logical states in the Schrödinger picture via EQ. 5 implies that (i) $\mathcal{W}$ teleports $k$ pairs of logical operators in the Heisenberg picture; (ii) $X^{(n)}_{{\rm L}}(T)\,=\,\mathcal{W}^{\dagger}\,X_{f_{n}}\,\mathcal{W}\,=\,X_{i_{n}}\,\mathsf{S}_{n,x}$ and $Z^{(n)}_{{\rm L}}(T)\,=\,\mathcal{W}^{\dagger}\,Z_{f_{n}}\,\mathcal{W}\,=\,Z_{i_{n}}\,\mathsf{S}_{n,z}$ are valid $X$- and $Z$-type logical operators for the $n$th logical qubit acting on the initial state $\lvert\Psi_{0}\rangle$ (6); and (iii) the operators $\mathsf{S}_{n,x}$ and $\mathsf{S}_{n,z}$ are elements of the initial-state stabilizer group $\operatorname*{stab}\left(*\right){\lvert\Psi_{0}\rangle}$ (3). Hence, Eq. 5 implies Eq. 1.
@@
Now, we prove the reverse implication: Any protocol $\mathcal{W}$ satisfying Eq. 1 with the initial state $\lvert\Psi_{0}\rangle$ (6) implies state transfer (5) with final state $\lvert\Psi_{T}\rangle$ (7)—i.e., Eq. 1 is sufficient for Eq. 5. For convenience, we define
$\rho_{n}\,$ $=\,\lvert\psi_{n}\rangle\langle\psi_{n}\rvert$
$=\,\frac{1}{2}\mathds{1}+c^{(n)}_{1}\,X_{i_{n}}++c^{(n)}_{2}\,Y_{i_{n}}+c^{(n)}_{3}\,Z_{i_{n}}$
@@
$c^{(n)}_{\nu}\,$ $\equiv\,\left\langle\psi_{n}\middle|\sigma^{\nu}_{i_{n}}\middle|\psi_{n}\right\rangle\,,~{}~{}$ (A.95)
@@
for the $n$th logical qubit, as in Eq. 2.2. We can then write the initial density matrix associated with $\lvert\Psi_{0}\rangle$ (6) as
@@
$\rho_{0}$ $=\lvert\Psi_{0}\left(\boldsymbol{\alpha},\boldsymbol{\beta}\right)\rangle\langle\Psi_{0}\left(\boldsymbol{\alpha},\boldsymbol{\beta}\right)\rvert\,$
@@
$=\left[\prod\limits_{n=1}^{k}\,\rho_{n}\right]\otimes\lvert\Phi_{i}\rangle\langle\Phi_{i}\rvert_{\rm rest}\,,$ (A.96)
@@
which realizes $\rho_{n}=\lvert\psi_{n}\rangle\langle\psi_{n}\rvert$ on each initial logical qubit $i_{n}$. Using Eqs. A.96 and A.94, we write
$\mathcal{W}\,\lvert\Psi_{i}\rangle$ $=\mathcal{W}\,\lvert\psi_{n}\rangle\langle\psi_{n}\rvert_{i_{n}}\lvert\Psi_{i}\rangle$
@@
$=\mathcal{W}\,\frac{1}{2}\left(\mathds{1}+\sum_{\nu=1}^{3}c^{(n)}_{\nu}\sigma^{\nu}_{i_{n}}\right)\lvert\Psi_{i}\rangle$
@@
$=\mathcal{W}\,\frac{1}{2}\left(\mathds{1}+\sum_{\nu=1}^{3}c^{(n)}_{\nu}\sigma^{\nu}_{i_{n}}\mathsf{S}_{n,\nu}\right)\lvert\Psi_{i}\rangle$
@@
$=\mathcal{W}\,\frac{1}{2}\left(\mathds{1}+\sum_{\nu=1}^{3}c^{(n)}_{\nu}\mathcal{W}^{\dagger}\sigma^{\nu}_{f_{n}}\,\mathcal{W}\right)\lvert\Psi_{i}\rangle$
@@
$=\frac{1}{2}\left(\mathds{1}+\sum_{\nu=1}^{3}c^{(n)}_{\nu}\sigma^{\nu}_{f_{n}}\right)\mathcal{W}\,\lvert\Psi_{i}\rangle$
$=\lvert\psi_{n}\rangle\langle\psi_{n}\rvert_{f_{n}}\,\mathcal{W}\,\lvert\Psi_{i}\rangle\,,~{}~{}$ (A.97)
@@
where the second line follows from Eq. A.94; the third line uses $\mathsf{S}_{n,\nu}\in\operatorname*{stab}\left(*\right){\lvert\Psi_{0}\rangle}$ to insert stabilizers inside the parenthetical expression; the fourth line uses Eqs. 1 and 9; and the last line uses Eq. A.94 once again. Thus, the final state $\lvert\Psi_{T}\rangle=\mathcal{W}\,\lvert\Psi_{0}\rangle$ (7) is stabilized by the $n$th logical qubit’s density matrix on site $f_{n}$. For brevity, we represent the final density matrix as $\rho_{T}=\mathcal{W}\,\lvert\Psi_{0}\rangle\langle\Psi_{0}\rvert\mathcal{W}^{\dagger}$. Since $\lvert\psi_{n}\rangle\langle\psi_{n}\rvert_{f_{n}}$ only acts nontrivially on site $f_{n}$, we have
11 11 $={\rm tr}\left[\,\rho_{T}\,\right]={\rm tr}\left[\,\lvert\psi_{n}\rangle\langle\psi_{n}\rvert_{f_{n}}\,\rho_{T}\,\right]$
$={\rm tr}\left[\,\lvert\psi_{n}\rangle\langle\psi_{n}\rvert\,\rho_{f_{n}}\,\right]=F\left(\rho_{f_{n}},\lvert\psi_{n}\rangle\right)\,,~{}$ (A.98)
@@
where $F$ is the fidelity Jozsa (1994), and in second line we used Eq. A; in the third line we have defined $\rho_{f_{n}}\equiv\operatorname*{tr}_{f^{c}_{n}}\left[\,\rho_{T}\,\right]$ as the reduced final density matrix on site $f_{n}$; in the last line we have used the fact that the fidelity obeys $F(\rho,\lvert\psi\rangle)=\left\langle\psi\middle|\rho\middle|\psi\right\rangle={\rm tr}\left[\,\lvert\psi\rangle\langle\psi\rvert\,\rho\,\right]$ when one of the states is pure Jozsa (1994). Note that, in the second line the trace is over the entire system, whereas in the third line, it is only over $f_{n}$. Since $F(\rho_{f_{n}},\lvert\psi_{n}\rangle)=1$, we must have that
@@
$\rho_{f_{n}}=\operatorname*{tr}_{f^{c}_{n}}\left[\,\rho\,\right]=\lvert\psi_{n}\rangle\langle\psi_{n}\rvert\,,~{}$ (A.99)
@@
i.e., the final reduced density matrix on site $f_{n}$ is the logical $n$th (pure) state. Since the above arguments hold for all $n=1,\dots,k$ independently, we conclude that $\lvert\Psi_{T}\rangle=\mathcal{W}\,\lvert\Psi_{0}\rangle$ satisfies (5) for the final state $\lvert\Psi_{T}\rangle$ (7). Thus, Eq. 1 is both necessary and sufficient for Eq. 5. ∎
@@ Appendix B Proof of Proposition 4
A generic and nontrivial single-qubit observable $A_{j}$ is guaranteed to have exactly two unique eigenvalues $a_{0}>a_{1}$, with spectral decomposition (15)
$A_{j}$ $=a_{0}\,P^{(0)}_{j}+a_{1}\,P^{(1)}_{j}\,,~{}~{}$ (B.100)
@@
$A_{j}$ $=\frac{1}{2}\,\sum\limits_{\nu=0}^{3}\,\operatorname*{tr}\limits_{j}\,\left[\,A_{j}\,\sigma^{\nu}_{j}\,\right]\,\sigma^{\nu}_{j}$
@@
where $\vec{\sigma}_{j}=(X_{j},Y_{j},Z_{j})$ is a vector of Pauli matrices and $\vec{\alpha}_{j}$ is a vector of coefficients as determined by the Pauli decomposition. Importantly, using the properties of $2\times 2$ matrices and the fact that $\det\,\left[A_{j}\right]=a_{0}\,a_{1}$ we find
$\lvert\vec{\alpha}\rvert=\frac{1}{2}\lvert a_{0}-a_{1}\rvert\,,~{}~{}$ (B.102)
@@
We rewrite the decomposed observable $A_{j}$ (B.101) as
$A_{j}$ $=\frac{1}{2}\left(a_{0}+a_{1}\right)\,\mathds{1}+\frac{1}{2}\left(a_{0}-a_{1}\right)\,\sum\limits_{n=1}^{3}\,\hat{\alpha}_{j,n}\,\sigma^{n}_{j}$
$=\frac{1}{2}\operatorname*{tr}\limits_{j}\left[A_{j}\right]\,\mathds{1}+\frac{1}{2}\,\left(a_{0}-a_{1}\right)\,\bar{A}_{j}\,,~{}~{}$ (B.103)
where the vector $\hat{\alpha}_{j}=(\hat{\alpha}_{j,1},\hat{\alpha}_{j,2},\hat{\alpha}_{j,3})$ is real valued and has unit length. This expression agrees with Eq. 4.
We next recover simple expressions for the projectors $P^{(n)}_{j}$ (B.100) onto the states (or spaces of states) with eigenvalues $a_{n}$. We first note that these two projectors must be orthogonal for the eigenvalue equation $A_{j}\,P^{(n)}_{j}=a_{n}\,P^{(n)}_{j}$ to hold. Second, we note that these two projectors must be complete (i.e., sum to the identity), or else evaluating $\det\,\left[A_{j}\right]$ in the eigenbasis of $A_{j}$ would return zero, independent of the values of $a_{0,1}$. As a result,
$A_{j}$ $=a_{n}\,P^{(n)}_{j}+a_{1-n}\left(\mathds{1}-P^{(n)}_{j}\right)$
@@
$=\frac{1}{a_{n}-a_{1-n}}\left(\frac{1}{2}\operatorname*{tr}\limits_{j}\left[A_{j}\right]\mathds{1}+\lvert\vec{\alpha}\rvert\,\bar{A}_{j}-a_{1-n}\mathds{1}\right)$
$=\frac{1}{a_{n}-a_{1-n}}\left(\frac{1}{2}(a_{n}-a_{1})\mathds{1}+\frac{1}{2}(a_{0}-a_{1})\,\bar{A}_{j}\right)$
$=\frac{1}{2}\left(\mathds{1}+(-1)^{n}\,\bar{A}_{j}\right)\,,~{}~{}$ (B.106)
where we used Eq. B.101 to write $A_{j}$ in terms of its involutory part $\bar{A}_{j}$ (4), along with the relations $a_{0}-a_{1}=(-1)^{n}(a_{n}-a_{1-n})=2\lvert\vec{\alpha}\rvert$ (B.102) and $\operatorname*{tr}_{j}\left[A_{j}\right]=a_{0}+a_{1}$.
Because the involutory operator $A_{j}$ squares to the identity, its eigenvalues are ±$\pm 1$. Accordingly, the projectors onto eigenstates of $A_{j}$ (which are also eigenstates of the original observable $A_{j}$) are given by Eq. B.106. Intuitively, these observables must share eigenstates because they commute: They differ only by a scalar factor and an overall, universally commuting identity component. Finally, because only the projectors onto eigenstates of $A_{j}$ appear in the unitary measurement channel (16), measurement of $A_{j}$ is equivalent to measurement of $\bar{A}_{j}$ (4). ∎

### v3

@@ Appendix A Proof of Proposition 1
An analogous procedure applies to the $Z^{\vphantom{*}}$-type logical operator for the $n$th logical qubit. That operator evolves in the Heisenberg picture according to
$Z^{(n)}_{{\rm L}}(T)\,=\,Z^{\vphantom{*}}_{f_{n}}(T)\,=\,\mathcal{W}^{\dagger}\,Z^{\vphantom{*}}_{f_{n}}\,\mathcal{W}\,,~{}~{}$ (A.92)
@@
in analogy to Eq. A.87, and implied by Eq. 7. We then define another stabilizer $\mathsf{S}^{\,}_{n,z}\equiv Z^{\vphantom{*}}_{i_{n}}\,\mathcal{W}^{\dagger}Z^{\vphantom{*}}_{f_{n}}\mathcal{W}$, implying that $\mathsf{S}^{\,}_{n,z}$ is also an element of $\operatorname*{stab}\left(*\right){\lvert\Psi^{\,}_{0}\rangle}$ (3). Applying the logical $Z^{\vphantom{*}}$ operator (A.92) to $\lvert\Psi^{\,}_{0}\rangle$ (6) reveals that
$Z^{(n)}_{{\rm L}}(T)=\mathcal{W}^{\dagger}\,Z^{\vphantom{*}}_{f_{n}}\,\mathcal{W}=Z^{\vphantom{*}}_{i_{n}}\,\mathsf{S}^{\,}_{n,z}\,,~{}~{}$ (A.93)
@@
is the $Z^{\vphantom{*}}$-type logical operator for the $n$th logical qubit acting on the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6), in analogy to Eq. A.91.
@@
Thus, the fact that $\mathcal{W}$ teleports $k$ logical states in the Schrödinger picture via EQ. 5 implies that (i) $\mathcal{W}$ teleports $k$ pairs of logical operators in the Heisenberg picture; (ii) $X^{(n)}_{{\rm L}}(T)\,=\,\mathcal{W}^{\dagger}\,X^{\vphantom{*}}_{f_{n}}\,\mathcal{W}\,=\,X^{\vphantom{*}}_{i_{n}}\,\mathsf{S}^{\vphantom{\prime}}_{n,x}$ and $Z^{(n)}_{{\rm L}}(T)\,=\,\mathcal{W}^{\dagger}\,Z^{\vphantom{*}}_{f_{n}}\,\mathcal{W}\,=\,Z^{\vphantom{*}}_{i_{n}}\,\mathsf{S}^{\vphantom{\prime}}_{n,z}$ are valid $X^{\vphantom{*}}$- and $Z^{\vphantom{*}}$-type logical operators for the $n$th logical qubit acting on the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6); and (iii) the operators $\mathsf{S}^{\,}_{n,x}$ and $\mathsf{S}^{\,}_{n,z}$ are elements of the initial-state stabilizer group $\operatorname*{stab}\left(*\right){\lvert\Psi^{\,}_{0}\rangle}$ (3). Hence, Eq. 5 implies Eq. 8.
@@
Now, we prove the reverse implication: Any protocol $\mathcal{W}$ satisfying Eq. 8 with the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6) implies state transfer (5) with final state $\lvert\Psi^{\,}_{T}\rangle$ (7)—i.e., Eq. 8 is sufficient for Eq. 5. For convenience, we define
$\rho^{\vphantom{\dagger}}_{n}\,$ $=\,\lvert\psi^{\,}_{n}\rangle\langle\psi^{\,}_{n}\rvert$
$=\,\frac{1}{2}\mathds{1}+c^{(n)}_{1}\,X^{\vphantom{*}}_{i_{n}}++c^{(n)}_{2}\,Y^{\vphantom{*}}_{i_{n}}+c^{(n)}_{3}\,Z^{\vphantom{*}}_{i_{n}}$
@@
$c^{(n)}_{\nu}\,$ $\equiv\,\left\langle\psi^{\,}_{n}\middle|\sigma^{\nu}_{i_{n}}\middle|\psi^{\,}_{n}\right\rangle\,,~{}~{}$ (A.95)
@@
for the $n$th logical qubit, as in Eq. 4. We can then write the initial density matrix associated with $\lvert\Psi^{\,}_{0}\rangle$ (6) as
@@
$\rho^{\,}_{0}$ $=\lvert\Psi^{\vphantom{\prime}}_{0}\left(\boldsymbol{\alpha},\boldsymbol{\beta}\right)\rangle\langle\Psi^{\vphantom{\prime}}_{0}\left(\boldsymbol{\alpha},\boldsymbol{\beta}\right)\rvert\,$
@@
$=\left[\prod\limits_{n=1}^{k}\,\rho^{\vphantom{\dagger}}_{n}\right]\otimes\lvert\Phi^{\vphantom{\prime}}_{i}\rangle\langle\Phi^{\vphantom{\prime}}_{i}\rvert^{\vphantom{\dagger}}_{\rm rest}\,,$ (A.96)
@@
which realizes $\rho^{\,}_{n}=\lvert\psi^{\,}_{n}\rangle\langle\psi^{\,}_{n}\rvert$ on each initial logical qubit $i_{n}$. Using Eqs. A.96 and A.94, we write
$\mathcal{W}\,\lvert\Psi^{\vphantom{*}}_{i}\rangle$ $=\mathcal{W}\,\lvert\psi^{\,}_{n}\rangle\langle\psi^{\,}_{n}\rvert^{\vphantom{*}}_{i_{n}}\lvert\Psi^{\vphantom{*}}_{i}\rangle$
@@
$=\mathcal{W}\,\frac{1}{2}\left(\mathds{1}+\sum_{\nu=1}^{3}c^{(n)}_{\nu}\sigma^{\nu}_{i_{n}}\right)\lvert\Psi^{\vphantom{*}}_{i}\rangle$
@@
$=\mathcal{W}\,\frac{1}{2}\left(\mathds{1}+\sum_{\nu=1}^{3}c^{(n)}_{\nu}\sigma^{\nu}_{i_{n}}\mathsf{S}^{\,}_{n,\nu}\right)\lvert\Psi^{\vphantom{*}}_{i}\rangle$
@@
$=\mathcal{W}\,\frac{1}{2}\left(\mathds{1}+\sum_{\nu=1}^{3}c^{(n)}_{\nu}\mathcal{W}^{\dagger}\sigma^{\nu}_{f_{n}}\,\mathcal{W}\right)\lvert\Psi^{\vphantom{*}}_{i}\rangle$
@@
$=\frac{1}{2}\left(\mathds{1}+\sum_{\nu=1}^{3}c^{(n)}_{\nu}\sigma^{\nu}_{f_{n}}\right)\mathcal{W}\,\lvert\Psi^{\vphantom{*}}_{i}\rangle$
$=\lvert\psi^{\,}_{n}\rangle\langle\psi^{\,}_{n}\rvert^{\vphantom{*}}_{f_{n}}\,\mathcal{W}\,\lvert\Psi^{\vphantom{*}}_{i}\rangle\,,~{}~{}$ (A.97)
@@
where the second line follows from Eq. A.94; the third line uses $\mathsf{S}^{\,}_{n,\nu}\in\operatorname*{stab}\left(*\right){\lvert\Psi^{\,}_{0}\rangle}$ to insert stabilizers inside the parenthetical expression; the fourth line uses Eqs. 8 and 9; and the last line uses Eq. A.94 once again. Thus, the final state $\lvert\Psi^{\,}_{T}\rangle=\mathcal{W}\,\lvert\Psi^{\,}_{0}\rangle$ (7) is stabilized by the $n$th logical qubit’s density matrix on site $f^{\,}_{n}$. For brevity, we represent the final density matrix as $\rho^{\,}_{T}=\mathcal{W}\,\lvert\Psi^{\,}_{0}\rangle\langle\Psi^{\,}_{0}\rvert\mathcal{W}^{\dagger}$. Since $\lvert\psi^{\,}_{n}\rangle\langle\psi^{\,}_{n}\rvert^{\,}_{f_{n}}$ only acts nontrivially on site $f_{n}$, we have
11 11 $={\rm tr}\left[\,\rho^{\vphantom{*}}_{T}\,\right]={\rm tr}\left[\,\lvert\psi^{\,}_{n}\rangle\langle\psi^{\,}_{n}\rvert^{\vphantom{*}}_{f_{n}}\,\rho^{\vphantom{*}}_{T}\,\right]$
$={\rm tr}\left[\,\lvert\psi^{\,}_{n}\rangle\langle\psi^{\,}_{n}\rvert\,\rho^{\vphantom{*}}_{f_{n}}\,\right]=F\left(\rho^{\vphantom{*}}_{f_{n}},\lvert\psi^{\vphantom{*}}_{n}\rangle\right)\,,~{}$ (A.98)
@@
where $F$ is the fidelity [78], and in second line we used Eq. A; in the third line we have defined $\rho^{\,}_{f_{n}}\equiv\operatorname*{tr}^{\,}_{f^{c}_{n}}\left[\,\rho^{\,}_{T}\,\right]$ as the reduced final density matrix on site $f_{n}$; in the last line we have used the fact that the fidelity obeys $F(\rho,\lvert\psi\rangle)=\left\langle\psi\middle|\rho\middle|\psi\right\rangle={\rm tr}\left[\,\lvert\psi\rangle\langle\psi\rvert\,\rho\,\right]$ when one of the states is pure [78]. Note that, in the second line the trace is over the entire system, whereas in the third line, it is only over $f_{n}$. Since $F(\rho^{\vphantom{*}}_{f_{n}},\lvert\psi^{\vphantom{*}}_{n}\rangle)=1$, we must have that
@@
$\rho^{\vphantom{*}}_{f_{n}}=\operatorname*{tr}^{\,}_{f^{c}_{n}}\left[\,\rho\,\right]=\lvert\psi^{\,}_{n}\rangle\langle\psi^{\,}_{n}\rvert\,,~{}$ (A.99)
@@
i.e., the final reduced density matrix on site $f_{n}$ is the logical $n$th (pure) state. Since the above arguments hold for all $n=1,\dots,k$ independently, we conclude that $\lvert\Psi^{\,}_{T}\rangle=\mathcal{W}\,\lvert\Psi^{\,}_{0}\rangle$ satisfies (5) for the final state $\lvert\Psi^{\,}_{T}\rangle$ (7). Thus, Eq. 8 is both necessary and sufficient for Eq. 5. ∎
@@ Appendix B Proof of Proposition 4
A generic and nontrivial single-qubit observable $A^{\,}_{j}$ is guaranteed to have exactly two unique eigenvalues $a^{\,}_{0}>a^{\,}_{1}$, with spectral decomposition (15)
$A^{\vphantom{\dagger}}_{j}$ $=a^{\,}_{0}\,P^{(0)}_{j}+a^{\,}_{1}\,P^{(1)}_{j}\,,~{}~{}$ (B.100)
@@
$A^{\vphantom{\prime}}_{j}$ $=\frac{1}{2}\,\sum\limits_{\nu=0}^{3}\,\operatorname*{tr}\limits_{j}\,\left[\,A^{\vphantom{\prime}}_{j}\,\sigma^{\nu}_{j}\,\right]\,\sigma^{\nu}_{j}$
@@
where $\vec{\sigma}^{\,}_{j}=(X^{\vphantom{*}}_{j},Y^{\vphantom{*}}_{j},Z^{\vphantom{*}}_{j})$ is a vector of Pauli matrices and $\vec{\alpha}^{\,}_{j}$ is a vector of coefficients as determined by the Pauli decomposition. Importantly, using the properties of $2\times 2$ matrices and the fact that $\det\,\left[A^{\,}_{j}\right]=a^{\,}_{0}\,a^{\,}_{1}$ we find
$\lvert\vec{\alpha}\rvert=\frac{1}{2}\lvert a^{\,}_{0}-a^{\,}_{1}\rvert\,,~{}~{}$ (B.102)
@@
We rewrite the decomposed observable $A^{\,}_{j}$ (B.101) as
$A^{\vphantom{\dagger}}_{j}$ $=\frac{1}{2}\left(a^{\,}_{0}+a^{\,}_{1}\right)\,\mathds{1}+\frac{1}{2}\left(a^{\,}_{0}-a^{\,}_{1}\right)\,\sum\limits_{n=1}^{3}\,\hat{\alpha}^{\vphantom{\prime}}_{j,n}\,\sigma^{n}_{j}$
$=\frac{1}{2}\operatorname*{tr}\limits_{j}\left[A^{\vphantom{\prime}}_{j}\right]\,\mathds{1}+\frac{1}{2}\,\left(a^{\,}_{0}-a^{\,}_{1}\right)\,\bar{A}^{\vphantom{\prime}}_{j}\,,~{}~{}$ (B.103)
where the vector $\hat{\alpha}^{\,}_{j}=(\hat{\alpha}^{\,}_{j,1},\hat{\alpha}^{\,}_{j,2},\hat{\alpha}^{\,}_{j,3})$ is real valued and has unit length. This expression agrees with Eq. 33.
We next recover simple expressions for the projectors $P^{(n)}_{j}$ (B.100) onto the states (or spaces of states) with eigenvalues $a^{\,}_{n}$. We first note that these two projectors must be orthogonal for the eigenvalue equation $A^{\,}_{j}\,P^{(n)}_{j}=a^{\,}_{n}\,P^{(n)}_{j}$ to hold. Second, we note that these two projectors must be complete (i.e., sum to the identity), or else evaluating $\det\,\left[A^{\,}_{j}\right]$ in the eigenbasis of $A^{\,}_{j}$ would return zero, independent of the values of $a^{\,}_{0,1}$. As a result,
$A^{\vphantom{\dagger}}_{j}$ $=a^{\,}_{n}\,P^{(n)}_{j}+a^{\,}_{1-n}\left(\mathds{1}-P^{(n)}_{j}\right)$
@@
$=\frac{1}{a^{\,}_{n}-a^{\,}_{1-n}}\left(\frac{1}{2}\operatorname*{tr}\limits_{j}\left[A^{\vphantom{\prime}}_{j}\right]\mathds{1}+\lvert\vec{\alpha}\rvert\,\bar{A}^{\vphantom{\prime}}_{j}-a^{\,}_{1-n}\mathds{1}\right)$
$=\frac{1}{a^{\,}_{n}-a^{\,}_{1-n}}\left(\frac{1}{2}(a^{\,}_{n}-a^{\,}_{1})\mathds{1}+\frac{1}{2}(a^{\,}_{0}-a^{\,}_{1})\,\bar{A}^{\vphantom{\prime}}_{j}\right)$
$=\frac{1}{2}\left(\mathds{1}+(-1)^{n}\,\bar{A}^{\vphantom{\prime}}_{j}\right)\,,~{}~{}$ (B.106)
where we used Eq. B.101 to write $A^{\,}_{j}$ in terms of its involutory part $\bar{A}^{\,}_{j}$ (33), along with the relations $a^{\,}_{0}-a^{\,}_{1}=(-1)^{n}(a^{\,}_{n}-a^{\,}_{1-n})=2\lvert\vec{\alpha}\rvert$ (B.102) and $\operatorname*{tr}_{j}\left[A^{\,}_{j}\right]=a^{\,}_{0}+a^{\,}_{1}$.
Because the involutory operator $A^{\,}_{j}$ squares to the identity, its eigenvalues are ±$\pm 1$. Accordingly, the projectors onto eigenstates of $A^{\,}_{j}$ (which are also eigenstates of the original observable $A^{\,}_{j}$) are given by Eq. B.106. Intuitively, these observables must share eigenstates because they commute: They differ only by a scalar factor and an overall, universally commuting identity component. Finally, because only the projectors onto eigenstates of $A^{\,}_{j}$ appear in the unitary measurement channel (16), measurement of $A^{\,}_{j}$ is equivalent to measurement of $\bar{A}^{\,}_{j}$ (33). ∎